Cornerstone Exercise 5.2 (Algorithmic) Job Costs Using a Plantwide Overhead Rate Naranjo Company designs industrial prototypes for outside companies. Budgeted overhead for the year was $290,000, and budgeted direct labor hours were 29,000. The average wage rate for direct labor is expected to be $20 per hour. During June, Naranjo Company worked on four jobs. Data relating to these four jobs follow: Job 39 Job 40 Job 41 Job 42 Beginning balance $26,700 $35,900 $16,700 $1,600 Materials requisitioned 20,300 23,200 12,600 14,200 Direct labor cost 11,400 20,300 7,250 5,100 Overhead is assigned as a percentage of direct labor cost. During June, Jobs 39 and 40 were completed; Job 39 was sold at 130 percent of cost. (Naranjo had originally developed Job 40 to order for a customer; however, that customer was near bankruptcy and the chance of Naranjo being paid was growing dimmer. Naranjo decided to hold Job 40 in inventory while the customer worked out its financial difficulties. Job 40 is the only job in Finished Goods Inventory.) Jobs 41 and 42 remain unfinished at the end of the month. Required: 1. Calculate the balance in Work in Process as of June 30. $ 2. Calculate the balance in Finished Goods as of June 30. $ 3. Calculate the cost of goods sold for June. $ 4. Calculate the price charged for Job 39. Round your answer to the nearest cent. $ 5. What if the customer for Job 40 was able to pay for the job by June 30? What would happen to the balance in Finished Goods? What would happen to the balance of Cost of Goods Sold?
Answer #1
job 39 job 40 job 41 job 42
balance 26700 35900 16700 1600
direct materials 20300 23200 12600 14200
direct labor 11400 20300 7250 5100
applied overhead 5700 10150 3625 2550
total 64100 89550 40175 23450

1. Calculate the balance in Work in Process as of June 30. = 40175+23450 = $63625

2. Calculate the balance in Finished Goods as of June 30. = $89550

3. Calculate the cost of goods sold for June. $64100

4. Calculate the price charged for Job 39. Round your answer to the nearest cent. $(64100*1.30) = $83330

5. What if the customer for Job 40 was able to pay for the job by June 30? What would happen to the balance in Finished Goods? Finished Goods would decrease

What would happen to the balance of Cost of Goods Sold?  Cost of Goods Sold would increase